OpenStudy (anonymous):

Write the equation y - 2 = 2(x - 3) in standard form.

OpenStudy (anonymous):

First, can you distribute the 2 on the right side to the parentheses?

OpenStudy (anonymous):

Ok.. y - 2 = (2x - 3)

OpenStudy (anonymous):

No, what is 2 times -3?

OpenStudy (ja1):

im not exactly spot on but im pretty sure its -2x - y = 4

OpenStudy (anonymous):

@JA1 please try to go through with the asker and explain how you got the answer, not just giving it.

OpenStudy (ja1):

well he asked "Write the equation y - 2 = 2(x - 3) in standard form."

OpenStudy (anonymous):

2 times -3 is -6

OpenStudy (ja1):

but i'll do that from now on

OpenStudy (anonymous):

@JA1 thanks! Ok @Umar97 , so the right side of the equation needs to be fixed now, the 2 was incorrectly distributed. Can you tell me what it is now?

OpenStudy (ja1):

no prob :)

OpenStudy (anonymous):

2x = 6

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

2x + 6*

OpenStudy (anonymous):

No, that is not. All you are working on is the right side right now:  Do you see how i got that?